# CapeNext: Rethinking and refining dynamic support information for category-agnostic pose estimation

### v1

@@ 1 Introduction
Figure 1: Conventional CAPE methods can not work when there is no support image but only textual keypoint descriptions, CapeX will be affected by the textual ambiguity to match the wrong candidate points, while our method can better avoid this problem.
@@ 3.2 Motivation
As illustrated in Figure 2, our framework CapeNext consists of three parts: (1) multimodal feature extraction and refinement, (2) feature matching and graph-based structural reasoning, and (3) keypoint prediction. Our key contribution lies in the design and refinement of multimodal feature which enhance support joint feature with both class-level and instance-specific cues. Transformer encoder and graph transformer decoder share the same architectures as GraphCape and CapeX.
@@ 3.3 Method Overview
Given a query image, we first extract its multi-scale visual features using a network backbone (e.g., HRNet-w32, ViT-Base-16, DINOv2-ViT-S and Swinv2-T). These features capture both local part details and global contextual patterns. To overcome polysemy-induced cross-category ambiguity and intra-category instance variation issues in CAPE, we derive an enhanced joint embedding through our multimodal input: textual class descriptions, which are easy to obtain, along with the query image itself as complementary information for keypoint and textual class descriptions.
@@ 4.1 Implementation Details
We perform experiments on MP-100 dataset (xu2022pose). The dataset consists of more than 18K images of 100 categories from 8 supercategories, including human hand, human face, human body, animal face, clothes, furniture, and vehicle. We also follow five splits for standard CAPE settings. Each split divides the dataset into non-overlap 70 training categories, 10 validation categories, and 20 testing categories. We also adopt the MP-100 dataset with unified skeleton definitions from GraphCape and keypoint text descriptions extended in CapeX. We primarily report probability of correct keypoint with a threshold of 0.2 (PCK@0.2) for evaluation. We also report the PCK performance of other thresholds (e.g.,0.05, 0.1, 0.15, 0.2, 0.25 and the avearge PCK performance of all thresholds) for comparison in our supplementary material.
@@ 4.2 Main Results
We compare CapeNext with eight SOTA methods including POMNet, CapeFormer, ESCAPE (nguyen2024escape), MetaPoint, X-Pose (yang2024x), SDPNet, GraphCape and our baseline CapeX. As Table 1 shows, with Swinv2-T as backbone, CapeNext outperforms other methods by an average of 0.76% under the 1-shot setting, validating the efficacy of our HCMI module and DSFR module. More results with HRNet-w32, ViT-Base-16 and DINOv2-ViT-S as backbone are also provided.
@@ 4.3 Ablation Study
Table 5: PCK@0.2 results when using different noisy keypoint prompts.

### v2

@@ 1 Introduction
Figure 1: Conventional CAPE methods can’t work without support image. With only textual keypoint descriptions, CapeX will be affected by the textual ambiguity to match the wrong candidate points, while our method works better.
@@ 3.2 Motivation
As illustrated in Figure 2, our framework CapeNext consists of three parts: (1) multimodal feature extraction and refinement, (2) feature matching and graph-based structural reasoning, and (3) keypoint prediction. Our key contribution lies in the design and refinement of multimodal feature which enhance support joint feature with both class-level and instance-specific cues. Transformer encoder and graph transformer decoder share the same architectures as CapeX.
@@ 3.3 Method Overview
Given a query image, we first extract its multi-scale visual features using a network backbone. These features capture both local part details and global contextual patterns. To overcome polysemy-induced cross-category ambiguity and intra-category instance variation issues in CAPE, we derive an enhanced joint embedding through our multimodal input: textual class descriptions, which are easy to obtain, along with the query image itself as complementary information for keypoint and textual class descriptions.
@@ 4.1 Implementation Details
We perform experiments on MP-100 dataset (xu2022pose). The dataset consists of more than 18K images of 100 categories, including human hand, human face, human body, animal face, clothes, furniture, and vehicle. We also follow five splits for standard CAPE settings. Each split divides the dataset into non-overlap 70 training categories, 10 validation categories, and 20 testing categories. We also adopt the MP-100 dataset with unified skeleton definitions from GraphCape and keypoint text descriptions extended in CapeX. We primarily report probability of correct keypoint with a threshold of 0.2 (PCK@0.2) for evaluation. We also report the PCK performance of other thresholds for comparison in our supplementary material.
@@ 4.2 Main Results
We compare CapeNext with eight SOTA methods including POMNet, CapeFormer, ESCAPE (nguyen2024escape), MetaPoint, X-Pose (yang2024x), SDPNet, GraphCape and our baseline CapeX. As Table 1 shows, with Swinv2-T as backbone, CapeNext outperforms other methods by an average of 0.76% under the 1-shot setting, validating the efficacy of our HCMI module and DSFR module. More results other backbones are also provided.
@@ 4.3 Ablation Study
Table 5: PCK@0.2 results when using noisy joint prompts.